OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9 A. (3, 9) B. (–27, –9) C. (–3, –6) D. (–9, –9)

OpenStudy (muzzack):

okay @Monster1234 @caleb_gajewsky

OpenStudy (anonymous):

mmmmmmmmmmm..............ahhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

when in doubt C your way out

OpenStudy (anonymous):

ikr

OpenStudy (muzzack):

3x – 4y = 9 (2)–3x + 2y = 9(2) --------- -6x+4y=18 --------- next:

OpenStudy (anonymous):

x=-9 right ?

OpenStudy (anonymous):

right I think so

OpenStudy (anonymous):

answer is (-9, -9)

OpenStudy (muzzack):

3x – 4y = 9 -6x+4y=18 9x=-9 x=-1

OpenStudy (anonymous):

and then its -9,-9

OpenStudy (muzzack):

did i make a mistake whatevs its D then lol

OpenStudy (anonymous):

hhha

OpenStudy (anonymous):

nobody's perfect <3